Operations Report – November 30th

Crew 216 Operations Report 30-Nov-19
SOL: 6
Name of person filing report: Mike Lawson

Non-nominal systems: RAM rover power charging stations inoperable, electrical issue, ongoing.
Notes on non-nominal systems:

Generator: run
Hours run: 11.5
From what time last night: 1902
To what time this morning: 0630
List any additional daytime hours when the generator was run: N/A

Solar— SOC % (Before the generator is run at night: 71%

Diesel Reading – 38
Station Propane Reading – 56
Ethanol Free Gasoline: N/A

Water (loft tank) (gal): 47
Water Meter (units): 1463316
Water (static tank) (gal): 315
Static to Loft Pump used – yes
Water in Green Hab (gal): 263
Water in Science Dome (gal): 0

Toilet tank emptied: yes

Deimos rover used: no
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used:
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: no
Hours: 114.5
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:

Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Curiosity rover used: no
Hours: 119.6
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100
Currently charging: yes

Notes on rovers: will have to juggle the charging sequence.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:

HabCar used and why, where?
CrewCar used and why, where?

General notes and comments: EVA to North Ridge, circumnavigated hill mass. No obvious routes for the summit in sim. No recon of potential radio repeater site.

Summary of internet: No serious issues.
Summary of suits and radios: No issues.
Summary of Hab operations: Routine maintenance and housekeeping.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: Gypsum processing continued.
Summary of RAM operations: RAM Rover power charging stations inoperable, electrical issues remain unresolved.
Summary of any observatory issues: Sun observations performed.